# Flexible and Probabilistic Topology Tracking with Partial Optimal Transport

### v3

@@ Abstract
publicationid: pubid: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 3.1 Merge Trees
The construction of a merge tree for a given $f:{\mathbb{M}}\to{\mathbb{R}}$ is described procedurally as follows: we sweep the function value $a$ from $\infty$- ∞ to $\infty$, and we create a new branch originating at a leaf node for each local minimum of $f$. As $a$ increases, such a branch is extended as its corresponding component in ${\mathbb{M}}_{a}$ grows until it merges with another branch at a saddle point. Assuming ${\mathbb{M}}$ is connected and $f$ achieves a unique global maximum, then all branches eventually merge into a single component, which corresponds to the root of the tree. For a given merge tree, leaves, internal nodes, and root node represent the minima, merging saddles, and global maximum of $f$, respectively. Fig. 1 displays a height function $f:{\mathbb{M}}\subset{\mathbb{R}}^{2}\to{\mathbb{R}}$ in (a), together with its corresponding merge tree embedded in the graph of the scalar field, i.e., $\{(x,f(x)):x\in{\mathbb{M}}\}$ in (b). Abstractly, a merge tree $T$ is a rooted tree equipped with $f$ restricted to its node set, $f:V\to{\mathbb{R}}$, as shown in (c).
@@ 3.3 Wasserstein and Gromov-Wasserstein Distance
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.4 Partial Wasserstein and Partial GW Distances
[…]
=$=\{C\in{\mathbb{R}}_{+}^{n_{1}\times n_{2}}\mid C{\mathbf{1}_{n_{2}}}\leq p_{1},C^{T}{\mathbf{1}_{n_{1}}}\leq p_{2},\mathbf{1}_{n_{1}}^{T}C{\mathbf{1}_{n_{2}}}=m\},$ (8)
@@
[…]
@@ 4.1 Partial Fused Gromov-Wasserstein Distance
For topology-based feature tracking, oftentimes features (i.e., critical points) will appear and disappear in time-varying data. Features that appear at time $t$ do not need to be matched with features at time $t-1$; similarly, features that disappear at time $t$ do not need to be matched with features at time $t+1$. Therefore, we need to introduce a partial Fused Gromov-Wasserstein (pFGW) distance for feature tracking to handle the appearances and disappearances of multiple features across time.
[…]
@@ 4.2.2 Encoding Extrinsic Information
[…]
@@
That is, the distance between categories is $1$ if the categories do not match and 00 if they do. This is referred to as the category strategy. In practice, we combine the above two distances to form the attribute distance, referred to as the combined strategy.
@@ 4.3.2 A Discussion on Flexibility
[…]
@@ 6 Experiments
We demonstrate the utility of our framework with five 2D datasets and two 3D datasets. For each dataset, we also compare against two state-of-the-art approaches. In particular, we demonstrate the strengths of our framework using two complex datasets—$\mathsf{Cloud}$ and $\mathsf{Viscous\,Fingering}$—in tracking a large number of features.
@@ 6.1 Datasets Overview
The $\mathsf{Heated\,Cylinder}$ dataset is a simulation of a 2D flow generated by a heated cylinder using the Boussinesq approximation [68, 69]. The simulation was done with a Gerris flow solver. It shows a time-varying turbulent plume containing numerous small vortices that, in part, rotate around each other. We generate a set of merge trees from the magnitude of the velocity fields based on $31$ time steps ($600$-$630$ from the original $2000$ time steps). These time steps describe the evolution of small vortices.
@@
The $\mathsf{Cloud}$ dataset shows the cloud optical thickness retrieved via the Daytime Cloud Optical and Microphysical Properties Algorithm (DCOMP) by Walther and Heidinger  [74], processed by Chatterjee et al. [75]. This 2D dataset has been used previously for cloud tracking [76]. We focus on the data sampled every 10 minutes from 10:50 to 16:50 on Feb 2, 2020 within the region of $10.82^{\circ}$ $15.88^{\circ}$, $49.19^{\circ}$ $42.51^{\circ}$. We use this dataset to demonstrate the utility of our pFGW framework on tracking a large number of features.
@@ 6.2.1 Parameter Tuning
Figure 5: $\mathsf{Heated\,Cylinder}$. Left: for a fixed $\alpha=0.1$, perform a grid search of $m$ and observe the number of oversegmentations (in blue boxes) and mismatches (in red boxes). Right: the trend among the number of trajectories ($N$) and the maximum matched distance ($L$) as $m$ increases.
@@ 6.5.1 Tracking Results
Figure 12: $\mathsf{Ionization\,Front}$. Left: pFGW trajectories are shown with the scalar field at time step 00. Right: pFGW trajectories are visualized with the landscape of the time-varying scalar field. Top: all trajectories; middle: long-term trajectories; bottom: short-term trajectories.
We demonstrate our pFGW tracking results in Fig. 12 (left), where trajectories are shown with the scalar field at $t=0$. We then visualize these trajectories with the landscape of the time-varying scalar field in Fig. 12 (right), which is constructed by stacking the original scalar field at time steps 00, $10$, $20$, …, $100$, and $110$. Such a landscape clearly shows the rightward propagation of the ionization front. The results shown in Fig. 12 (top) thus contain a number of trajectories that capture such a trend.
@@ 7 Probabilistic Tracking Graphs
We now visually demonstrate the probabilistic tracking graph for the $\mathsf{Ionization\,Front}$ dataset. In the example shown in Fig. 21, we set the probability threshold at $0.023$ (main view) and $0.032$ (inserted view), respectively. To investigate the data of interest, users could select a specific time step $t$ by clicking its corresponding time bar, which updates views (a), (b), and (c). For the graph view (a), the selected time bar is highlighted in red, with the previous two (at $t-2$, $t-1$) and subsequent two (at $t+1$, $t+2$) time bars colored in orange and blue, respectively. We smoothly adjust intervals between time steps based on the fisheye technique using animation, so that the focus area surrounding the selected time bar is magnified and the area away from the focus is compressed. For the track view (b), we render five scalar fields centered by the selected time step and highlight the tracks among them in a 3D spacetime, while supporting zooming and rotation. For the data view (c), we visualize the five 2D scalar fields side by side, where tracked features are highlighted in red.
@@ Appendix A A New Stability Theorem for GW Distance
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The claim is proved (setting $C=\frac{1}{4}$) if we can choose $2p_{0}^{2}=p_{1}$ such that $12n\cdot p_{0}+p_{1}=$ 1 and $p_{0},p_{1}>0$, so that $p$ defines a valid probability density. Solving the system of equations with the given constraints, we get
@@
[…]
@@ Appendix C Parameter Tuning Details
For the $\mathsf{Isabel}$ dataset: we did not include GFT results; LWM sets $\gamma=1$ and $\beta=1$. For pFGW, we use $\epsilon=10\%$, $\alpha=0.6$, $L^{*}=0.4010$. $L^{*}$ value is large because there are large gaps between time instances in this dataset and the hurricane also makes large movement across these time steps.
@@ D.2 Quantitative Comparisons
TABLE II: Similarity measures between a pair of tracking results with and without subsampling. For a fixed tracking method, $A$ denotes the trajectories restricted to subsampled time steps. $B$ denotes the trajectories from the subsampled dataset. The highest scores are in bold.

### v4

# largesymbols”00 largesymbols”01

@@ 3.1 Merge Trees
The construction of a merge tree for a given $f:{\mathbb{M}}\to{\mathbb{R}}$ is described procedurally as follows: we sweep the function value $a$ from -$\infty$ to $\infty$, and we create a new branch originating at a leaf node for each local minimum of $f$. As $a$ increases, such a branch is extended as its corresponding component in ${\mathbb{M}}_{a}$ grows until it merges with another branch at a saddle point. Assuming $\mathbb{M}$} is connected and $f$ achieves a unique global maximum, then all branches eventually merge into a single component, which corresponds to the root of the tree. For a given merge tree, leaves, internal nodes, and root node represent the minima, merging saddles, and global maximum of $f$, respectively. Fig. 1 displays a height function $f:{\mathbb{M}}\subset{\mathbb{R}}^{2}\to{\mathbb{R}}$ in (a), together with its corresponding merge tree embedded in the graph of the scalar field, i.e., $\{(x,f(x)):x\in{\mathbb{M}}\}$ in (b). Abstractly, a merge tree $T$ is a rooted tree equipped with $f$ restricted to its node set, $f:V\to{\mathbb{R}}$, as shown in (c).
@@ 3.3 Wasserstein and Gromov-Wasserstein Distance
We refer to $d_{A}(a_{i},b_{j})$ as the attribute distance between nodes $x_{i}\in V_{1}$ and $y_{j}\in V_{2}$. The Wasserstein distance aims to minimize the weighted sum of attribute distance between matched nodes. The minimizers in Eq. (2) are referred to as optimal couplings.
@@
FGW distance. Vayer et al. introduced the fused Gromov-Wasserstein (FGW) distance between attributed graphs and other structured objects [54, 21]. We describe their framework in the setting of measure networks. The FGW distance is a trade-off between the Wasserstein distance in Eq. (2) and the GW distance in Eq. (3). For $q\in[1,\infty)$ and a trade-off parameter $\alpha\in[0,1]$, the FGW distance between attributed measure networks $G_{1}$ and $G_{2}$ is defined (following [54]) as
@@
Here, $C$ is considered as a soft assignment matrix, and $\alpha$ gives a trade-off between labels and structures. As shown in Sec. 4, Eq. (4) plays an important role in encoding both intrinsic and extrinsic information for merge tree comparisons.
@@
As $\alpha\to 1$, the FGW distance recovers the GW distance (ignoring the constant factor in Eq. (3)),
@@ 3.4 Partial Wasserstein and Partial GW Distances
${\mbox{$\mathcal{C}$}}_{m}={\mathcal{C}}_{m}(p_{1},p_{2})$
$=\{C\in{\mathbb{R}}_{+}^{n_{1}\times n_{2}}\mid C{\mathbf{1}_{n_{2}}}\leq p_{1},C^{T}{\mathbf{1}_{n_{1}}}\leq p_{2},\mathbf{1}_{n_{1}}^{T}C{\mathbf{1}_{n_{2}}}=m\},$ (8)
@@
A main difference between partial Wasserstein distance and Wasserstein distance is that we replace the equalities in Eq. (1) with inequalities in Eq. (3.4) to account for “partial mass transport”.
@@ 4.1 Partial Fused Gromov-Wasserstein Distance
For topology-based feature tracking, oftentimes features (i.e., critical points) will appear and disappear in time-varying data. Features that appear at time $t$ do not need to be matched with features at time t1t-1; similarly, features that disappear at time $t$ do not need to be matched with features at time $t+1$. Therefore, we need to introduce a partial Fused Gromov-Wasserstein (pFGW) distance for feature tracking to handle the appearances and disappearances of multiple features across time.
The pFGW distance is defined based on the set of admissible couplings ${\mathcal{C}}_{m}$ in Eq. (3.4) and the FGW distance in Eq. (4). Given a pair of measure networks $G_{1}$ and $G_{2}$, formally, we have
@@ 4.2.2 Encoding Extrinsic Information
Extrinsic information that typically arises from the geometry of the data domain may be encoded via the attribute space $(A,d_{A})$ and attribute distance $d_{A}$ in Eq. (4). For a node in the merge tree, the assigned attribute may be a high-dimensional vector or a categorical label. Given a pair of merge trees $T_{1}=(V_{1},p_{1},W_{1})$ and $T_{2}=(V_{2},p_{2},W_{2})$, nodes $x_{i}\in V_{1}$ and $y_{j}\in V_{2}$ are equipped with attributes $a_{i}$ and $b_{j}$ from the same attribute space $(A,d_{A})$.
@@
That is, the distance between categories is $1$ if the categories do not match and 0 if they do. This is referred to as the category strategy. In practice, we combine the above two distances to form the attribute distance, referred to as the combined strategy.
@@ 4.3.2 A Discussion on Flexibility
Second, parameter $\alpha$ from Eq. (11) strikes a balance in considering intrinsic information (via the GW distance) and extrinsic information (via the Wasserstein distance) for merge tree comparisons.
Third, parameter $m$ from Eq. (11) allows partial mass transport to accommodate the appearances and disappearances of features.
@@ 6 Experiments
[…]
@@ 6.1 Datasets Overview
The $\mathsf{Heated\,Cylinder}$ dataset is a simulation of a 2D flow generated by a heated cylinder using the Boussinesq approximation [68, 69]. The simulation was done with a Gerris flow solver. It shows a time-varying turbulent plume containing numerous small vortices that, in part, rotate around each other. We generate a set of merge trees from the magnitude of the velocity fields based on $31$ time steps (600600-$630$ from the original $2000$ time steps). These time steps describe the evolution of small vortices.
@@
The $\mathsf{Cloud}$ dataset shows the cloud optical thickness retrieved via the Daytime Cloud Optical and Microphysical Properties Algorithm (DCOMP) by Walther and Heidinger  [74], processed by Chatterjee et al. [75]. This 2D dataset has been used previously for cloud tracking [76]. We focus on the data sampled every 10 minutes from 10:50 to 16:50 on Feb 2, 2020 within the region of $10.82^{\circ}$N - $15.88^{\circ}$N, $49.19^{\circ}$W - $42.51^{\circ}$W. We use this dataset to demonstrate the utility of our pFGW framework on tracking a large number of features.
@@ 6.2.1 Parameter Tuning
Figure 5: $\mathsf{Heated\,Cylinder}$. Left: for a fixed $\alpha=0.1$, perform a grid search of $m$ and observe the number of oversegmentations (in blue boxes) and mismatches (in red boxes). Right: the trend among the number of trajectories (NN) and the maximum matched distance ($L$) as $m$ increases.
@@ 6.5.1 Tracking Results
Figure 12: $\mathsf{Ionization\,Front}$. Left: pFGW trajectories are shown with the scalar field at time step 0. Right: pFGW trajectories are visualized with the landscape of the time-varying scalar field. Top: all trajectories; middle: long-term trajectories; bottom: short-term trajectories.
We demonstrate our pFGW tracking results in Fig. 12 (left), where trajectories are shown with the scalar field at $t=0$. We then visualize these trajectories with the landscape of the time-varying scalar field in Fig. 12 (right), which is constructed by stacking the original scalar field at time steps 0, $10$, 2020, …, $100$, and $110$. Such a landscape clearly shows the rightward propagation of the ionization front. The results shown in Fig. 12 (top) thus contain a number of trajectories that capture such a trend.
@@ 7 Probabilistic Tracking Graphs
We now visually demonstrate the probabilistic tracking graph for the $\mathsf{Ionization\,Front}$ dataset. In the example shown in Fig. 21, we set the probability threshold at $0.023$ (main view) and $0.032$ (inserted view), respectively. To investigate the data of interest, users could select a specific time step $t$ by clicking its corresponding time bar, which updates views (a), (b), and (c). For the graph view (a), the selected time bar is highlighted in red, with the previous two (at t2t-2, t1t-1) and subsequent two (at $t+1$, $t+2$) time bars colored in orange and blue, respectively. We smoothly adjust intervals between time steps based on the fisheye technique using animation, so that the focus area surrounding the selected time bar is magnified and the area away from the focus is compressed. For the track view (b), we render five scalar fields centered by the selected time step and highlight the tracks among them in a 3D spacetime, while supporting zooming and rotation. For the data view (c), we visualize the five 2D scalar fields side by side, where tracked features are highlighted in red.
@@ Appendix A A New Stability Theorem for GW Distance
The right-hand side of Eq. (14) uses a weighted $L^{q}$ norm as defined above, for functions on $V\times V$. That is, for $F:V\times V\to\mathbb{R}$, we have
@@
To establish Eq. (16), we refer to Lemma 1. Let $v,w\in V$ be arbitrary vertices and assume without loss of generality that $W_{f}(v,w)\geq W_{g}(v,w)$. We have
@@
Finally, Eq. (17) follows from a general bound on $q$-norms:
@@
Next, the proof appears to be inefficient in that no optimization over couplings is used and that we pass through an $\infty$-norm rather than working directly with $q$-norms throughout. However, Eq. (13) is asymptotically tight up to a universal constant, as is shown by the following example. Let $V=\{v_{1},\ldots,v_{2n+1}\}$ and form $X$ by joining $v_{i}$ to $v_{i+1}$ by an edge for $i=1,\ldots,2n$. Define functions $f,g:V\to\mathbb{R}$ by setting $g(v_{i})=0$ for all $i$ and
@@
The claim is proved (setting $C=\frac{1}{4}$) if we can choose $2p_{0}^{2}=p_{1}$ such that $2n\cdot p_{0}+p_{1}=1$ and $p_{0},p_{1}>0$, so that $p$ defines a valid probability density. Solving the system of equations with the given constraints, we get
@@
and the claim follows, once we justify the steps. Eq. (19) holds by the same reasoning as in the proof of Theorem 2, Eq. (20) is an application of Eq. (18), Eq. (21) is the triangle inequality for the $L^{q}$ norm, together with marginalizing out a copy of $p$ in the second term, and Eq. (22) follows from the proof of Theorem 2. ∎
@@ Appendix C Parameter Tuning Details
For the $\mathsf{Isabel}$ dataset: we did not include GFT results; LWM sets $\gamma=1$ and $\beta=1$. For pFGW, we use $\epsilon=10\%$, $\alpha=0.6$, $^{*}=0.4010. LL^{*}$ value is large because there are large gaps between time instances in this dataset and the hurricane also makes large movement across these time steps.
@@ D.2 Quantitative Comparisons
TABLE II: Similarity measures between a pair of tracking results with and without subsampling. For a fixed tracking method, AA denotes the trajectories restricted to subsampled time steps. $B$ denotes the trajectories from the subsampled dataset. The highest scores are in bold.